# Particle motion around luminous neutron stars: effects of deviation from Schwarzschild spacetime

### v1

@@ IV Discussion and summary
We demonstrated the importance of the spacetime geometry for equilibria and dynamics of particles around luminous compact stars. In the astrophysical context, understanding the location of the gravity-radiation equilibrium surface is important for the observational constraints on the neutron star equation of state from the photospheric radius expansion X-ray bursts in the touchdown phase Ozel2006 ; Galloway2008 ; Kim2021 , or from oscillations of the material suspended above the star surface by the strong radiation Bollimpalli2019 . On the one hand, relaxing the Schwarzschid spacetime model assumption results in another source of uncertainty, on the other hand, it may offer a way to test strong gravity. What is more, the dynamical properties of such systems could manifest observationally, e.g., increased velocity of the expanding atmosphere of a bursting neutron star would necessarily manifest as a shift in the observed X-ray spectrum. These problems have not received sufficient attention from the scientific community so far.
There are multiple research avenues that should be followed in order to generalize the presented results and make them more physically relevant. In particular, the test-particle approximation can be relaxed and more realistic cases of optically thin Wielgus2015b ; Bollimpalli2017 and optically thick Wielgus2016 levitating (suspended in the radiation field) atmospheres can be considered. The questions regarding the impact of the rotation of the star and the role of magnetic fields are also important as subjects of further studies that will necessarily render the presented toy model more complicated.
@@ Acknowledgements
We thank Marek Abramowicz and Włodek Kluźniak for discussions. We also thank Alexandra Elbakyan for her contributions to the open science initiative. MW acknowledges support from the European Research Council advanced grant “M2FINDERS - Mapping Magnetic Fields with INterferometry Down to Event hoRizon Scales” (Grant No. 101018682).

### v2

@@ IV Discussion and summary
We demonstrated the importance of the spacetime geometry for equilibria and dynamics of particles around luminous compact stars. In the presented examples of spherically symmetric spacetimes deviating from the Schwarzschild solution, the radius of the equilibrium sphere can be larger by as much as a factor of 2 with respect to the Schwarzschild equilibrium at the same intrinsic luminosity. Apart from this effect constituting an interesting theoretical problem, there are potentially important applications of the presented results in the astrophysical context of phenomena involving strong radiation originating in the vicinity of a compact gravitating object. Understanding the location of the gravity-radiation equilibrium surface is relevant for the observational constraints on the neutron star equation of state from the type-I luminous thermonuclear photospheric radius expansion X-ray bursts Ozel2006 ; Galloway2008 ; Kim2021 , or from oscillations of the material suspended above the star surface by the strong radiation Bollimpalli2019 . In the former case, as an example, the ratio between observed near-Eddington fluxes at the peak of the X-ray burst and near its end (the touchdown phase) provides information on the gravitational redshift at the stellar surface Damen1990 , related to our parameter $\xi_{R}$. The relation between the inferred redshift and the $M/R$ quantity that can be subsequently used to constrain the neutron star equation of state is thus depending on the underlying spacetime geometry through the metric function $\xi(r)$. Similarly, for the measurement proposed in Bollimpalli2019 , the frequencies of the equilibrium sphere oscillations depend on the spacetime geometry and the related location of the equilibrium, affecting the inferred mass and radius of the neutron star.
On the one hand, relaxing the Schwarzschild spacetime model assumption results in another source of uncertainty, on the other hand, it may offer a way to test strong gravity. What is more, the dynamical properties of such systems could manifest observationally, e.g., increased velocity of the expanding atmosphere of a bursting neutron star would necessarily manifest as a shift in the observed X-ray spectrum. These problems have not received sufficient attention from the scientific community so far.
There are multiple research avenues that should be followed in order to generalize the presented results and make them more physically relevant. In particular, the test-particle approximation can be relaxed and more realistic cases of optically thin Wielgus2015b ; Bollimpalli2017 and optically thick Wielgus2016 levitating (suspended in the radiation field) atmospheres can be considered. The questions regarding the impact of the rotation of the star and the role of magnetic fields are also important as subjects of further studies that will necessarily render the presented toy model more complicated. In particular, (Wielgus2019, ) demonstrated that for a rotating source with a realistic radiation stress-energy tensor (Miller1996, ) the radial equilibrium is only present in the star’s equatorial plane, with radiation drag forces operating efficiently in the poloidal direction, destabilizing off-equatorial equilibria on relevant timescales for dimensionless spins as small as $a=0.05$. For a fixed luminosity and other system parameters, the equatorial equilibrium occurs at a slightly larger radius if the star is rotating. Thus, a toroidal configuration, puffed-up by the pressure of the accumulated gas around the test-particle equatorial equilibrium is expected. Such two-dimensional equilibria have not been calculated so far (see (Wielgus2015b, ; Wielgus2016, ) for a non-rotating spherically symmetric case).
@@ Acknowledgements
We thank Marek Abramowicz and Włodek Kluźniak for discussions. We also thank Alexandra Elbakyan for her contributions to the open science initiative.